# GMG: A Video Prediction Method Based on Global Focus and Motion Guided

### v2

@@ IV-C Qualitative Evaluation
Fig. 6 presents the visualization results for WeatherBench and Moving MNIST. We highlight the main error-prone regions in WeatherBench, and for better observation, all errors are magnified tenfold. GMG achieves the smallest error among all models. Although the variations in this dataset are relatively slower compared to rainfall data, the combined effect of the Global Focus Module and Motion Guided Module still enables the model to capture detailed temperature changes effectively.
For the Moving MNIST dataset, we provide visualizations comparing GMG with TAU. At the final prediction step (T=10), TAU produces a relatively blurry prediction of the digit ”8” in the image, while GMG generates a much clearer result that is closer to the ground truth. Although the Motion Guided Module is primarily designed for non-rigid motion features, it also performs well in rigid motion prediction tasks. This demonstrates that GMG is versatile, adaptable to a wide range of scenarios, and has the potential to extend to more datasets.
[…]
From the above visualization results, we observe that GMG can adapt to various types of video prediction tasks, regardless of whether the predicted subject is a rigid or elastic body. Moreover, it achieves superior performance in terms of prediction accuracy and object detail preservation. This demonstrates that GMG can serve as a general-purpose video prediction model, making it applicable to real-world complex video prediction tasks.
IV-D Ablation Study
To validate the contributions of GMG’s individual components, we conduct ablation experiments on the CIKM2017 dataset, with results presented in Table VI. In addition to controlled experiments on key components—GFM, MGM, and SAM—we include comparisons with baseline models, PredRNN and MotionRNN. The results show that adding GFM, SAM, or MGM improves predictive performance over the baselines. Furthermore, compared to MotionGRU, model (A), which incorporates the MGM module to account for growth or dissipation processes, better captures rainfall data characteristics. By comparing Model (A), (B), and (C), MGM emerges as the most impactful component in enhancing baseline performance, followed by GFM, further validating the effectiveness of the proposed model design. Additionally, the lightweight version, GMG-s, achieves competitive results, offering a viable pathway for future lightweight modeling efforts. We also provide the parameter count and training time for each model, offering a reference for practical model deployment.
TABLE VI: Ablation study results on CIKM2017 dataset.
Model index GFM SAM MGM Params Traning Time (per epoch) MSE (↓) MAE (↓) SSIM (↑)
PredRNN 23.8M 380s 28.8126 159.404 0.7303
MotionRNN MotionGRU 26.9M 425s 27.2091 155.827 0.7406
(c) 26.9M 440s 26.4196 153.801 0.7462
(d) 31.4M 516s 26.9540 155.525 0.7444
(e) 30.4M 538s 26.7479 155.658 0.7448
(f) 31.4M 627s 25.2383 151.723 0.7528
GMG-s (Ours) Simplified 29.2M 541s 25.3616 151.334 0.7501
GMG-m (Ours) 30.4M 455s 26.1861 152.472 0.7487
GMG-L (Ours) 31.44M 589s 25.0215 149.493 0.7513
@@ IV-D Ablation Study
Extensive experiments on five benchmark datasets demonstrate that GMG significantly enhances prediction performance, surpassing existing mainstream methods while exhibiting strong generalization and application potential across different tasks.

### v3

@@ IV-D Ablation Study
ConvLSTM 40.0678 16.1892 0.9819
+ GFM 35.2095 ($\Delta$-4.8583) 15.2149 ($\Delta$-0.9743) 0.9842 ($\Delta$+0.0023)
PredRNN 35.0348 15.1302 0.9844
+ GFM 32.8259 ($\Delta$-2.2089) 14.9933 ($\Delta$-0.1369) 0.9846 ($\Delta$+0.0002)
PredRNNpp 41.8227 15.9766 0.9824
+ GFM 32.3835 ($\Delta$-9.4392) 15.2241 ($\Delta$-0.7525) 0.9842 ($\Delta$+0.0018)
MotionRNN 37.6517 16.0009 0.9825
+ GFM 34.9205 ($\Delta$-2.7312) 15.2291 ($\Delta$-0.7718) 0.9838 ($\Delta$+0.0013)
PredRNNv2 45.2737 16.6105 0.9807
+ GFM 39.0525 ($\Delta$-6.2212) 15.3422 ($\Delta$-1.2683) 0.9833 ($\Delta$+0.0026)
TABLE XII: Comparison of MotionRNN with MotionGRU and Motion Guided Module (MGM) across three datasets.
Dataset Model MSE (↓) MAE (↓) SSIM (↑)
CIKM2017 MotionRNN 27.2091 155.827 0.7406
+MGM 26.4060 ($\Delta$-0.8031) 154.574 ($\Delta$-1.253) 0.7445 ($\Delta$+0.0039)
Shanghai2020 MotionRNN 4.5867 37.871 0.9221
+MGM 4.4530 ($\Delta$-0.1337) 36.971 ($\Delta$-0.900) 0.9246 ($\Delta$+0.0025)
TaxiBJ MotionRNN 0.3765 16.000 0.9825
+MGM 0.3408 ($\Delta$-0.0357) 15.297 ($\Delta$-0.703) 0.9840 ($\Delta$+0.0015)
@@
Extensive experiments on six benchmark datasets demonstrate that GMG significantly enhances prediction performance, surpassing existing mainstream methods while exhibiting strong generalization and application potential across different tasks.